OpenStudy (anonymous):

Simplify the roots. Click on my question please to see the problem. It is a pic.

OpenStudy (anonymous):

\[(\sqrt{63}) + (\sqrt{7})=\]

OpenStudy (tkhunny):

Hint: 63 = 9*7